Find the ordered pair (x,y) that satisfies the system of equations
x + y = 15
x - y = 7 + 4x - 19y

Found 2 solutions by amarjeeth123, ikleyn:
Answer by amarjeeth123(569)

The equations are as given below.
x + y = 15
x - y = 7 + 4x - 19y
These equations can be rewritten as given below.
x+y=15
-3x+18y=7
3x+3y=45
-3x+18y=7
21y=52
y=52/21
x=15-(52/21)=263/21.

Answer by ikleyn(52790)

.

Let me tell you that this formulation in the post

        "Find the ordered pair (x,y) that satisfies the system of equations"

is too wordy for Math and looks strange.


        Word  " ordered "  is excessive.

        Word  " pair "  is excessive.

        Word  " satisfies "  is excessive.


Too many unnecessary words make writing unprofessional.

Much better is to say "Solve this system of equations".


            Simple and convincing.